OpenStudy (anonymous):

A zoo train ride costs $5 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who took the ride was 17, and the total money collected was $65. What was the number of children and the number of adults who took the train ride that day, and which pair of equations can be solved to find the numbers?

OpenStudy (anonymous):

A. 5 childrenand 12 adults Equation 1: a + c = 17 Equation 2: 5a – c = 65 B. 12 childrenand 5 adults Equation 1: a + c = 17 Equation 2: 5a + c = 65 C. 12 childrenand 5 adults Equation 1: a + c = 17 Equation 2: 5a – c = 65 D. 5 childrenand 12 adults Equation 1: a + c = 17 Equation 2: 5a + c = 65

OpenStudy (anonymous):

those are the answers!

OpenStudy (anonymous):

b

OpenStudy (anonymous):

so its right?

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

@welshfella is that right?

OpenStudy (welshfella):

a + c = 17 5a + c = 65 - these equations fit the numbers 12 + 5 = 17 5*12 + 5 = 65

OpenStudy (welshfella):

b

OpenStudy (anonymous):

k

OpenStudy (welshfella):

OH NOOO sorry its d because its 5 dollars for adult and 1 for child D is correct

OpenStudy (anonymous):

wow so you changed it? are you sure its right?

OpenStudy (welshfella):

a is the number of adults and c = number of children

OpenStudy (welshfella):

yes I'm sure

OpenStudy (anonymous):

?